# Hovering Flight in Flapping Insects and Hummingbirds: A Natural Real-Time and Stable Extremum Seeking Feedback System

### v3

@@ Abstract
In this paper, we take an initial and novel step towards characterizing the physical phenomenon of hovering flight as an extremum seeking (ES) feedback system. We anticipate that said novel characterization may start a new line of research that can potentially solve all the puzzle pieces of hovering flight that existed for decades in previous literature. Is hovering flight stable? If so, what is the control mechanism utilized by insects/hummingbirds to achieve stable hovering? If such a mechanism exists, does it fit the biological constraints that insects/hummingbirds have limited computational abilities? Does it fit the experimental biology narrative that insects/hummingbirds rely mainly on their sensation to stabilize hovering? Our ES characterization and analysis provide for the first time a simple, model-free, real-time, stable feedback system of hovering. Consistent with natural observations and biological experiments, hovering via ES is simply achievable by the natural oscillations of the wing angle and measuring (sensing) altitude or acceleration. We provide simulation trials, including comparisons with some approaches from literature, to demonstrate the effectiveness of our results. We used literature data for hawkmoth, cranefly, bumblebee, dragonfly, hoverfly, and a hummingbird.
@@ I Introduction
Flapping flight is a fascinating physical phenomenon that captivated the interests of different communities such as physicists, biologists, and aerospace/control engineers. For decades, flapping flight has been a rich research area with a highly complex nature, involving multi-body, nonlinear, and non-autonomous dynamics [1, 2]. Moreover, the mechanics of flapping-wing flight impose a unique challenge on researchers in many fields, as conventional aerodynamic principles developed for fixed-wing aircraft are quite limited in characterizing flapping-wing systems [1, 3]. This is especially true when it comes to the flapping flight of insects and hummingbirds due to the ability of said organisms to hover.
Hovering flight is a unique, puzzling physical phenomenon for biologists and engineers. It was once deemed impossible under traditional aerodynamic theory, as the required hovering lift coefficients for balance are significantly higher than what is achievable through conventional aerodynamics [4, 5, 6, 7, 8, 1, 9]. Over time, unconventional lift mechanisms that can physically enable hovering flight has been discovered, particularly the concept of the leading-edge vortex, introduced in [10] and later expanded upon by the authors in [11], which provides the necessary lift for insect hovering flight.
Consequently, significant progress has been made in both theoretical and experimental research to better understand, analyze, and mimic hovering flight [1, 2, 3]. On the theoretical side, researchers tried to derive and advance the modeling aspects of hovering flight (on the body and aerodynamic levels) to analyze properties such as stability. They also aimed to introduce control mechanisms/techniques that can explain stable hovering and enable the application of hovering in robotics. On the experimental side, most efforts are aimed at studying flapping insects and the relation between their sensation and the ability to hover. Nevertheless, the gap between theoretical studies (more on that in Subsections I.1 and I.2) in literature and experimental works (Subsection I.3) is wide and irreconcilable, motivating the need for the contributions of this paper (Subsection I.4).
I.1 Is hovering flight stable or not?
The long-held view and consensus in the literature is that hovering flapping flight is inherently unstable and active control may be required for stabilization [1, 9, 12, 13, 14, 15, 16]. A recent study [17] extended stability analysis to a broader range of hovering insects has reached the same conclusion as well. In all aforementioned studies and in many others [18, 19, 12, 20, 21, 22, 23, 24, 25, 26], stability analysis was conducted based on two main approximations. The first is a physical approximation, where the wing’s inertial effects are neglected relative to body dynamics due to its smaller mass (typically less than 5% [27]), reducing the system to a single-body problem. The second is a mathematical approximation, where direct averaging [28] (also known as first-order averaging [29]) is applied to analyze balance and stability by converting the nonlinear time-periodic hovering system into a nonlinear time-invariant (autonomous) averaged system. Equipped by the two approximations mentioned above, various kinds/modes of dynamics have also been considered in stability analysis, such as but not limited to 3-degree of freedom (3DOF) longitudinal flight system or the full 6-degree of freedom (6DOF) system [25, 27, 30, 31]. Moreover, several studies have used computational fluid dynamics (CFD) models to simulate the aerodynamic forces acting on insects/ hummingbirds during 3DOF and 6DOF hovering flight [21, 22, 27]. Once again, all mentioned studies suggested that hovering flapping flight is unstable and may need active control. It should be noted that the stability-instability debate in the aforementioned literature is usually regarding the pitching motion of flapping insects.
Recently, the authors in [32, 33] argued that concluding instability of hovering in literature is due to a deficiency in the mathematical approximation (i.e., first-order averaging) and that hovering flight may be shown stable, including with pitching motion, in an open-loop sense if higher-order averaging is used [32]. However, they were able to counter the literature results (i.e., demonstrate stability) only for the case of hawkmoth. Therefore, they acknowledged that feedback control may be required for other insects to stabilize their hovering flight. Said acknowledgment [33], as well as the literature consensus about the need for active control for stable hovering, aligns with the work of the authors in [34] who discerned the feedback control mechanism used by fruit flies to maintain stable flight dynamics; more on feedback in insects will be discussed in Subsection I.3.
I.2 Hovering flight control in literature
Due to the consensus in literature (as discussed in previous Subsection I.1) that hovering flight is inherently unstable, researchers then tried to introduce control techniques/mechanisms that may enable stabilized hovering flight [1, 9, 35]. However, it is important to highlight that control techniques used for hovering in literature rely on model-dependent approaches at the design level and/or the implementation level, requiring knowledge of the equilibrium hovering point (i.e., the objective function) a priori. In addition, in some works, aerodynamic models necessary for the control function were based on CFD simulations, which are inherently non-real-time. As such, very simple, real-time, model-free control for hovering in flapping insects cannot be found in literature prior to this paper, to the best of our knowledge. Now, we briefly mention some of the control mechanisms that have been used in hovering flight literature. Broadly speaking, since controlling the hovering flight in literature is model-dependent, researchers often have to determine two main characteristics. First, the dynamic and aerodynamic models they use to represent the insect/mimicking-system. Most control studies for hovering flapping flight utilize either the standard aircraft equations of motion or custom multi-body models designed to capture the complex interactions between the wings and the body [36, 37, 38, 37, 38]. In [39, 24] a simplified blade-element/quasi-steady aerodynamic model was used based on the work of [40, 41], and the authors in [21, 42, 43] used more detailed CFD models to simulate the forces generated during flapping. These simulations were used to determine stability derivatives and the effects of various control inputs in hover and forward flight. They concluded that CFD models provide higher accuracy, especially for capturing unsteady aerodynamic effects, but they are computationally expensive. Second, the control design/structure itself which actuates stable hovering. Standard designs such as Linear Quadratic Regulator (LQR) (e.g., [44]) and Proportional-Integral-Derivative (PID) (e.g., [16, 45]) exists in literature. Additionally, the reader can refer to some other nonlinear control designs such as but not limited to [37, 38, 46].
As noted in literature, model-based control designs have their downsides since they inherently depend on the accuracy of the dynamic/aerodynamic models. Increasing their accuracy is quite challenging computationally; we leave the reader with this quotation from [47] which expands on this point:“Although, at present, some numerical simulations of unsteady insect flight aerodynamics based on the finite element solution of the Navier–Stokes equations give accurate results for the estimated aerodynamics forces [48, 49], their implementation is unsuitable for control purposes since they require several hours of processing for simulating a single wingbeat, even on multiprocessor computers.”
I.3 Natural and Biological Observations
Natural and biological observations/experiments are in consensus that flapping insects rely heavily on their sensation, which means it is highly plausible that they use some feedback control mechanism to stabilize their hovering in real-time based on their sensations – see for example [34, 12, 50, 51, 52]. It is worth highlighting that there is considerable literature that adopt the opinion, and argue, that flapping insects have very limited computational abilities – see for example relevant discussions on this in [33, 53]. As a result, researchers tried via experiments to investigate how then insects rely on their sensation to perform their highly efficient flapping flight, including hovering. Now we mention a few relevant experiments and observations from literature. The authors in [35, 34] examined how freely flying fruit flies react in real-time to external disturbances affecting their yaw and pitch. Small ferromagnetic materials were attached to the flies, allowing for reorientation via a magnetic field. Observations (via three high-speed cameras) indicated that the flies adapted their wing kinematics in real-time in a manner similar to that used in voluntary maneuvers, implying the utilization of real-time active auto-stabilization control mechanisms. Other studies as well have explored the sensory mechanisms that flapping insects utilize for navigation and stability. For instance, an experimental study by [54] demonstrated that flapping insects rely on a combination of optical systems (compound eyes and ocelli), airflow sensors (antennae and wind-sensitive hairs), and inertial sensors (halteres) to maintain stability during flight. In a subsequent experiment, they deactivated the halteres (insect gyroscopes), leaving the fly reliant solely on visual feedback, which responds approximately four times slower. As a result, the fly was unable to sustain flight, suggesting that intrinsic flight stability is absent without haltere input (i.e., without sufficient feedback). Interestingly, stability was restored when light dandelion fibers were attached to the fly’s abdomen, indicating the importance of sensory feedback in stabilization. Furthermore, [55, 56] highlighted the role of the ocellar system in various insect species. The ocelli, consisting of three broad-field photoreceptors located on the heads of many flying insects, function as crucial sensory components. Positioned to capture light from different areas of the sky, these receptors help insects maintain stable flight, suggesting that the ocelli act as horizon detectors by evaluating changes in light intensity across the photoreceptors. Consequently, based on the aforementioned studies and as noted by others [33, 47, 12, 50, 51, 57], it is strongly believed that flapping insects utilize sensory feedback to achieve a controlled stable hovering.
I.4 Motivation and Contribution
Figure 1: The literature on hovering flight in flapping insects/hummingbirds has irreconcilable contradictions as conclusions about instability when pitching motion is included and proposed feedback control methods are at odds with biological observations and experiments.
The findings of the literature in previous Subsections (summarized in Figure 1) are contradictory in many ways, deeming hovering flight as an unresolved puzzle. For instance, conclusions about the instability of hovering contradict our own eyes and natural observations that flapping insects and hummingbirds indeed stabilize their hovering for certain objectives (e.g., reaching a food source at a particular altitude). On the other hand, proposing feedback control methods in literature as active control means to achieve stable hovering has fallen short. That is, available control methods in literature, while possessing the feedback feature highlighted by biological experiments are model-dependent. That is, they require knowledge of the equilibrium hovering point (i.e., the objective function) a priori. Additonally, some control methods in literature can be computationally expensive and challenging to perform in real-time.
In this paper, we take an initial, and novel, step towards characterizing the hovering phenomenon in flapping insects and hummingbirds as an a new class of extremum seeking (ES) feedback systems [58, 59, 60]. Said novel characterization, while in early phases, strikes a great leap towards solving all the puzzle pieces of hovering flight that existed for decades in previous literature as depicted in Figure 1. In fact, during the last decade, the use of ES feedback systems have achieved breakthroughs in characterizing/mimicking important and challenging behaviors in biological organisms immersed in fluid structures in a model-free fashion; see for example ES in fish [61] and some microswimmers [62, 63]. Very recently, our group, for the first time in literature, introduced ES characterization in flying organisms and succeeded in resolving the decades-long problem of decoding/mimicking the energy-efficient flight strategy of soaring birds [64, 65].
@@ I.4 Motivation and Contribution
By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing (i.e., the natural flapping action) as the ES input, and feedback of sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the contradictions highlighted in Figure 1. It also may provide model-free, real-time, computationally cheap, bio-mimicry possibilities for micro-air-vehicles. In this paper, we test our proposed hovering ES (see Section III) using literature data for five flapping insects (hawkmoth, cranefly, bumblebee, dragonfly, hoverfly) and a hummingbird using simplified models (see Section II) representing the body dynamics and aerodynamics of flapping insects/hummingbirds. Said testing is for the proof of concept as ES systems work in a model-free fashion [58, 59, 60]. We also compare our ES approach with open-loop setting and a PID control approach. Testing and simulation results, including with feedback noise and delay, indicate success in stabilizing hovering for the aforementioned insects and hummingbird based on literature data. Following that (in Section IV) we also provide stability analysis using sophisticated tools from differential geometric control theory [66, 67]. Stability results indicate stable hovering for all tested insects and hummingbird. Finally, we conclude the paper in Section V.
@@ II Modelling of Flapping Insects and Hummingbirds: Body Dynamics and Aerodynamics
Now recall from Subsection I.1, two primary approximations have been considered in flapping flight modeling, namely physical approximations and mathematical approximations. We only revisit the physical approximations since they can affect the accuracy of the model itself (note that mathematical approximations are merely for stability analysis, so they do not affect the modeling accuracy when simulating the flapping system). Recently, there have been some developments [32, 70, 71, 72, 73, 68, 69, 74] that we think are worth consideration in our modeling. In [73], the authors demonstrated that the interaction between wing flapping dynamics and body flight dynamics can lead to a negative lift mechanism, significantly affecting overall flight performance. Additionally, on the aerodynamic side, unsteady aerodynamics models [70, 71, 72] for flapping have gained traction as they include factors like leading-edge vortex, rotational dynamics, viscous and added mass effects, and wake capture, all of which contribute to a more accurate representation of aerodynamic loads.
@@ III Natural Hovering flight as an extremum seeking system
Extremum seeking (ES) systems [58, 59, 60] are themselves control systems (called sometimes ESC). They steer/drive a dynamical system to the extremum point (maximum/minimum) of an objective function that can be unknown, expression-wise. ES systems have been powerful and attractive for many reasons, including that they are (1) very simple structurally and in implementation, (2) stable, (3) model-free, and (4) real-time. In fact ES systems have found many applications in real-time dynamic optimization and control systems in many fields (for more details, the reader can refer to [60]). With relevance to the topic of this paper, and as mentioned in Subsection I.4, ES systems have had significant successes in characterizing and mimicking some optimized behaviors in biological organisms such as fish [61], micorswimmers [62, 63], and soaring birds [64, 65].
@@
The classic ES system [77, 58] in its simplest forms (the one that is structurally closer to what we use in this paper) can be considered as an input-output system. The input is a perturbation (modulation) action/signal, which causes variation in the system dynamics. This translates into variations (corresponding to input perturbations/modulation) in the objective function, which is considered the output in ES systems and the only feedback needed. Unlike most control techniques, in ES systems, even the objective function, including its maximum/minimum point, can be unknown. All that the ES system needs is access to its measurements. Unfamiliar readers with why this model-free nature of ES (including dealing with unknown objective functions) is particularly useful for the paper at hand, and is a near-unique property among control techniques, may refer to these two papers regarding model-free, real-time light-source seeking [78, 79] by ES systems – see also this video [80]. In these papers (and the video), the light source, which represents the maximum of the objective function, that is, the light signal itself, is unknown position-wise; the ES system steers the robot to the light source without any information about the position of the source, but by simply measuring (sensing) light. In Subsection III.C, we will see how seeking unknown position/altitude is much more challenging using control methods such as PID. In this paper, we characterize and treat hovering flight as an ES system (i.e., ES is the model and control simultaneously for hovering). First, we expand on the philosophy/reasoning behind our hypothesis before moving into our attempt to provide a proof of concept.
@@
Hovering insects/hummingbirds [1, 2, 3] utilize flapping as their action. To be more precise, and to expand the word “flapping” in physical terms, they conduct very high oscillatory motion of their flapping wing angle $\phi$. In kinematic terms, this is usually translated into a high amplitude sinusoidal function. For the flapping model we use in this paper (see (8)), said oscillatory motion is associated with the torque $\tau$, which is responsible for changing the flapping angle rate $\dot{\phi}$. Thus, perturbations in $\tau$ is the modulation step. Due to perturbation/modulation in $\tau$, $\dot{\phi}$ is updated, causing variation in $\phi$. As a consequence, the dynamic system (including the insect’s body) will be varied as well, resulting in variation in any measurements (sensation by the insect) that depend on the state variables of the system dynamics. Let us clarify this further using the light source seeking analogy we discussed earlier (including citations to these papers [78, 79] and this video [80]). Let us assume that a flapping insect (e.g., a moth) is stabilizing its altitude (i.e., hovering) based on the sensation of light emitting from a source with unknown position. This means that the altitude about which the insect will stabilize corresponds to the unknown maximum light intensity. If we successfully characterize hovering flight as an ES system, this implies that via measuring (sensing) altitude or altitude-related signals, the insect will be able to stabilize about the desired altitude, which corresponds to the maximum light intensity. Moreover, in source seeking applications, such as light source seeking, a perturbation signal is typically introduced into the control input as an integral part of the ES control system [78, 79]. This perturbation enables the system to sense the objective function and adjust accordingly. However, in the natural hovering ES system, this additional perturbation is unnecessary, as it is inherently present due to the wing flapping motion itself. The oscillatory nature of flapping wings naturally induces this perturbation, allowing the system to sense the objective function in real-time without any external modifications, even when the objective function is unknown. Based on the biological observations in Subsection I.3 and the light source seeking discussion above, we rationally assume that the insect/hummingbird is able to sense their altitude (or altitude-related signals). Hence, without loss of generality, we assume the objective function $J$ to reflect altitude difference between current state and desired altitude to stabilize about certain altitude. The hovering ES system aims to minimize $J$, so $J$ measurements (sensation) is demodulated, and is used to update the torque average (mean estimate) $\hat{\tau}$. Then, another cycle starts by perturbation again. Our natural hovering ES system is provided in Figure 4.
@@
One major difference between our proposed natural hovering ES system in Figure 4 and the traditional classic ES structure in [77, 58] is the nature of the perturbation/modulation (also the demodulation) signals. In traditional classic ES systems, the amplitude of the modulation and demodulation signals is considered small (e.g., $a\>cos(\Omega t)$ for relatively small $a$). However, in the proposed natural hovering ES system, we use high amplitude, high frequency perturbation signals in the form $a\>\Omega cos(\Omega t$) following the natural/biological observations in literature which are reflected in the flapping model we use in (8). Hence, the perturbation input, flapping angle torque $\tau$, is given by $\tau=\hat{\tau}+a\Omega\cos(\Omega t)$. The objective function $J$ which represents the output $J=g(x)$ will be either $J=z^{2}$ representing altitude or altitude-related signals, or $J=\dot{w}^{2}$ representing acceleration or acceleration related signals as shown in Figure 4. This sets up the ES system to steer the insect towards the minimum of $J$ by updating the flapping angle based on updating $\tau$. It is important to note that in classic ES systems [77, 58] high and/or low pass filters can be used optionally. In our natural hovering ES system (Figure 4), no filters at all are used. Now, we are in a position to provide the proposed natural hovering ES system as in (9) and (10):

### v4

@@ Abstract
In this paper, we take an initial and novel step toward characterizing the hovering phenomenon in flapping insects and hummingbirds as a new class of extremum seeking (ES) feedback systems. By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing as the ES input. That is, unlike other control techniques in the literature, the natural hovering ES system only needs the natural flapping action built in the system, and feedback of local sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the long-standing gap between model-based control theory and the biologically observed mechanisms that stabilize hovering flight. We provide simulation trials, including comparisons with some approaches from literature, to demonstrate the effectiveness and robustness of our results. We used literature data for hawkmoth, cranefly, bumblebee, dragonfly, hoverfly, and a hummingbird.
@@ I Introduction
Flapping flight is a fascinating physical phenomenon that captivated the interests of different communities such as physicists, biologists, and aerospace/control engineers. For decades, flapping flight has been a rich research area with a highly complex nature, involving multi-body, nonlinear, and non-autonomous dynamics [1, 2]. Moreover, the mechanics of flapping-wing flight impose a unique challenge on researchers in many fields, as conventional aerodynamic principles developed for fixed-wing aircraft are quite limited in characterizing flapping-wing systems [1, 3]. This is especially true when it comes to the flapping flight of insects and hummingbirds due to the ability of said organisms to hover. Hovering flight is a unique, puzzling physical phenomenon for biologists and engineers. It was once deemed impossible under traditional aerodynamic theory, as the required hovering lift coefficients for balance are significantly higher than what is achievable through conventional aerodynamics [4, 5, 6, 7, 8, 1, 9]. Over time, unconventional lift mechanisms that can physically enable hovering flight has been discovered, particularly the concept of the leading-edge vortex, introduced in [10] and later expanded upon by the authors in [11], which provides the necessary lift for insect hovering flight. Consequently, significant progress has been made in both theoretical and experimental research to better understand, analyze, and mimic hovering flight [1, 2, 3]. On the theoretical side, researchers tried to derive and advance the modeling aspects of hovering flight (on the body and aerodynamic levels) to analyze properties such as stability. They also aimed to introduce control mechanisms/techniques that can explain stable hovering and enable the application of hovering in robotics. On the experimental side, most efforts are aimed at studying flapping insects and the relation between their sensation and the ability to hover.
The long-held view and consensus in the literature is that hovering flapping flight is inherently unstable in the open-loop sense and active control may be required for stabilization [1, 9, 12, 13, 14, 15, 16, 17]. A recent study [18] extended stability analysis to a broader range of hovering insects has reached the same conclusion as well. Recently, the authors in [19, 20] argued that concluding instability is due to a deficiency in first-order averaging and that hovering flight may be shown stable—including the pitching motion in an open-loop sense if higher-order averaging is used; however, they acknowledged that feedback control may be required for other insects to stabilize their hovering flight. Natural and biological observations/experiments are in consensus that flapping insects rely heavily on their sensation, which means it is highly plausible that they use some feedback control mechanism to stabilize their hovering in real time [21, 22, 12, 17, 23, 24]. The authors in [17, 22] examined how freely flying fruit flies react in real time to external disturbances affecting their yaw and pitch: small ferromagnetic materials were attached to the flies, allowing for reorientation via a magnetic field, and observations indicated that the flies adapted their wing kinematics in real-time. Thus, this can imply that, in the absence of passive damping, active sensorimotor feedback is important for stabilization.
From a control-theoretic perspective, stabilized hovering is achieved through sensory feedback [20, 25, 12, 17, 26, 21]. This understanding has motivated many researchers to develop feedback-based control techniques aimed at enabling stable hovering. Consequently, a variety of model-dependent control strategies have been proposed in the literature [1, 9]. Standard approaches such as Linear Quadratic Regulator (LQR), Proportional-Integral-Derivative (PID) control, and other nonlinear designs rely on dynamic and aerodynamic models that require knowledge of the equilibrium hovering point a priori [27, 28, 29, 30]. In several studies, aerodynamic models necessary for control were based on computational fluid dynamics simulations, which are inherently non-real-time [31, 28, 32, 33]. It is worth mentioning that, although PID and its variants (PI and PD) are real-time controllers, even the PI scheme that reproduces yaw recovery in fruit flies [17] — as the authors themselves state — requires prior estimation of morphological and aerodynamic parameters and subsequent hand-tuning of gain parameters, with all parameters varying by ±$\pm 15\%$. Moreover, none of the control methods, techniques, or mechanisms discussed above leverage the inherent high-frequency wing-flapping perturbation action, which can serve as a built-in probing signal.
Motivation. The consensus that hovering flight is open-loop unstable and needs feedback control for stabilization has motivated researchers to introduce a variety of feedback control techniques to achieve stable hovering; however, available control methods can be model-dependent, require knowledge of the equilibrium hovering point a priori, and can potentially be computationally expensive to implement in real-time due to tuning challenges and/or high need of computational power. Moreover, none of these controllers leverage the inherent high-frequency wing-flapping perturbation action built into the system. Hence, we are motivated to propose a control method that can frame the hovering problem while using the built-in high-frequency wing perturbation action in a natural way as the probing signal for the controller itself. For that, we turn our attention to Extremum Seeking (ES) control systems [34, 35, 36], which utilize perturbation actions as the probing signal for the controller to derive a given dynamical system towards the extremum of unknown objective functions in a model-free, real-time, sensation-based fashion. Moreover, ES feedback control systems have achieved breakthroughs in characterizing and mimicking challenging behaviors in biological organisms immersed in fluid structures — see, for example, fish [37] and microswimmers [38, 39]. Recently, our group introduced ES characterization in flying organisms and succeeded in decoding the energy-efficient soaring strategy of birds [40, 41].
@@
Contribution. In this paper, we take an initial and novel step toward characterizing the hovering phenomenon in flapping insects and hummingbirds as a new class of ES feedback systems. By characterizing hovering flight in insects and hummingbirds as a natural hovering ES system, we achieve: (1) very simple, (2) stable, (3) model-free, and (4) real-time hovering. More importantly, our hovering ES characterization only needs the natural oscillations of the wing (i.e., the natural flapping action, built in the system, unlike other control techniques in the literature) as the ES input, and feedback of local sensations (measurements) related to the altitude where the insect seeks to stabilize itself. Said ES characterization, can become an important initial step in starting a new line of research that may succeed in resolving the long-standing gap between model-based control theory and the biologically observed mechanisms that stabilize hovering flight. It also may provide model-free, real-time, computationally cheap, bio-mimicry possibilities for flapping robots.
In this paper, we test our proposed hovering ES (see Section III) using literature data for five flapping insects (hawkmoth, cranefly, bumblebee, dragonfly, hoverfly) and a hummingbird using simplified models (see Section II) representing the body dynamics and aerodynamics of flapping insects/hummingbirds. Said testing is for the proof of concept as ES systems work in a model-free fashion [34, 35, 36]. We also compare our ES approach with open-loop setting and a PID control approache, including optimal-gains PID controllers. Testing and simulation results, including with feedback noise and delay, indicate success in stabilizing hovering for the aforementioned insects and hummingbird based on literature data. Following that (in Section IV) we also provide stability analysis using sophisticated tools from differential geometric control theory [42, 43, 44]. Stability results indicate stable hovering for all tested insects and hummingbird. Finally, we conclude the paper in Section V.
@@ II Modelling of Flapping Insects and Hummingbirds: Body Dynamics and Aerodynamics
There are two primary approximations that have been considered in flapping flight modeling, namely physical approximations and mathematical approximations. We only revisit the physical approximations since they can affect the accuracy of the model itself (note that mathematical approximations are merely for stability analysis, so they do not affect the modeling accuracy when simulating the flapping system). Recently, there have been some developments [19, 47, 48, 49, 50, 45, 46, 51] that we think are worth consideration in our modeling. In [50], the authors demonstrated that the interaction between wing flapping dynamics and body flight dynamics can lead to a negative lift mechanism, significantly affecting overall flight performance. Additionally, on the aerodynamic side, unsteady aerodynamics models [47, 48, 49] for flapping have gained traction as they include factors like leading-edge vortex, rotational dynamics, viscous and added mass effects, and wake capture, all of which contribute to a more accurate representation of aerodynamic loads.
@@
Remark. It is to be noted that the model in (8) is a reduced and simplified model for hovering. Nevertheless, as long as this reduced model is relatively accurate in capturing the essential qualitative features of hovering, it is sufficient for testing the ES framework. This is due to the fact that ES systems are inherently model–free and tolerant of large model uncertainties and/or variations [36]; in our case, this will be further illustrated in Section III.D. That is, successful performance and stabilization in hovering for the simplified model adopted in this work, suggest that the same ES framework will remain effective under the richer, more variable dynamics of real insects. In other words, the ES framework is robust to modeling errors.
@@ III Natural Hovering flight as an extremum seeking system
Extremum seeking (ES) systems [34, 35, 36] are themselves control systems (called sometimes ESC). They steer/drive a dynamical system to the extremum point (maximum/minimum) of an objective function that can be unknown, expression-wise. ES systems have been powerful and attractive for many reasons, including that they are (1) very simple structurally and in implementation, (2) stable, (3) model-free, and (4) real-time. In fact ES systems have found many applications in real-time dynamic optimization and control systems in many fields (for more details, the reader can refer to [36]). With relevance to the topic of this paper, ES systems have had significant successes in characterizing and mimicking some optimized behaviors in biological organisms such as fish [37], micorswimmers [38, 39], and soaring birds [40, 41].
@@
ES systems [55, 34, 36], generally speaking, can be considered as input-output systems. The input is a perturbation (modulation) action/signal, which causes variation in the system dynamics. This translates into variations (corresponding to input perturbations/modulation) in the objective function, which is considered the output in ES systems and the only feedback required. Recently, inspired by the flapping mechanics of insects, a class of ES systems was proposed [44], which is relevant to systems possessing mechanical structure (e.g., systems described by Newton’s second law or Euler-Lagrange formulation) with generalized forces that are quadratic in velocities; this matches, and can be applied to, the system representing hovering at hand in (8). Unlike most control techniques, in ES systems, even the objective function, including its maximum/minimum point, can be unknown. All that the ES system needs is access to its local measurements. Unfamiliar readers with why this model-free nature of ES (including dealing with unknown objective functions) is particularly useful for the paper at hand, and is a near-unique property among control techniques, may refer to these two papers regarding model-free, real-time light-source seeking [56, 57] by ES systems – see also this video [58]. In these papers (and the video), the light source, which represents the maximum of the objective function, that is, the light signal itself, is unknown position-wise; the ES system steers the robot to the light source without any information about the position of the source, but by simply measuring (sensing) light. In this paper, we characterize and treat hovering flight as an ES system (i.e., ES is the model and control simultaneously for hovering). First, we expand on the philosophy/reasoning behind our hypothesis before moving into our attempt to provide a proof of concept.
@@
Hovering insects/hummingbirds [1, 2, 3] utilize flapping as their action. To be more precise, and to expand the word “flapping” in physical terms, they conduct very high oscillatory motion of their flapping wing angle $\phi$. In kinematic terms, this is usually translated into a high amplitude sinusoidal function. For the flapping model we use in this paper (see (8)), said oscillatory motion is associated with the torque $\tau$, which is responsible for changing the flapping angle rate $\dot{\phi}$. Thus, perturbations in $\tau$ is the modulation step. Due to perturbation/modulation in $\tau$, $\dot{\phi}$ is updated, causing variation in $\phi$. As a consequence, the dynamic system (including the insect’s body) will be varied as well, resulting in variation in any measurements (sensation by the insect) that depend on the state variables of the system dynamics. Let us clarify this further using the light source seeking analogy we discussed earlier (including citations to these papers [56, 57] and this video [58]). Let us assume that a flapping insect (e.g., a moth) is stabilizing its altitude (i.e., hovering) based on the sensation of light emitting from a source with unknown position. This means that the altitude about which the insect will stabilize corresponds to the unknown maximum light intensity. If we successfully characterize hovering flight as an ES system, this implies that via local measuring (sensing) altitude or altitude-related signals, the insect will be able to stabilize about the desired altitude, which corresponds to the maximum light intensity. Moreover, in source seeking applications, such as light source seeking, a perturbation signal is typically introduced into the control input as an integral part of the ES control system [56, 57]. This perturbation enables the system to sense the objective function and adjust accordingly. However, in the natural hovering ES system, this additional perturbation is unnecessary, as it is inherently present due to the wing flapping motion itself. The oscillatory nature of flapping wings naturally induces this perturbation, allowing the system to sense the objective function in real-time without any external modifications, even when the objective function is unknown.
It is worth to mention and emphasize that the light–seeking illustration is only one example of how local sensation or measurements might be enough to seek an extremum. But this local sensation (measurements) could be heat, light, chemical concentration, among many other possibilities. For example, the authors in [39] apply ES to sperm that cannot “see” the egg and rely only on chemotaxis.
Based on the biological observations and the light source seeking discussion above, we rationally assume that the insect/hummingbird is able to sense their altitude (or altitude-related signals). Hence, without loss of generality, we assume the objective function $J$ to reflect the altitude difference between the current state and the desired altitude to stabilize about certain altitude. The hovering ES system aims to minimize $J$, so $J$ measurements (sensation) is demodulated, and is used to update the torque average (mean estimate) $\hat{\tau}$. Then, another cycle starts by perturbation again. Our natural hovering ES system is provided in Figure 3.
@@
One major difference between our proposed natural hovering ES system in Figure 3 and the traditional classic ES structure in [55, 34] is the nature of the perturbation/modulation (also the demodulation) signals. In traditional classic ES systems, the amplitude of the modulation and demodulation signals is considered small (e.g., $a\>cos(\Omega t)$ for relatively small $a$). However, in the proposed natural hovering ES system, we use high amplitude, high frequency perturbation signals in the form $a\>\Omega cos(\Omega t$) following the natural/biological observations in literature, which are reflected in the flapping model we use in (8). As a result, our proposed natural ES system is a member of the ES classes in [44]. Furthermore, the perturbation input, flapping angle torque $\tau$, is given by $\tau=\hat{\tau}+a\Omega\cos(\Omega t)$. The objective function $J$ which represents the output $J=g(x)$ will be either $J=z^{2}$ representing altitude or altitude-related signals, or $J=\dot{w}^{2}$ representing acceleration or acceleration related signals as shown in Figure 3. This sets up the ES system to steer the insect towards the minimum of $J$ by updating the flapping angle based on updating $\tau$. Now, we are in a position to provide the proposed natural hovering ES system as in (9) and (10):